### v3

@@ 3.2.2 Digital Twins for 6G Network native AI
In addition, DT channel is also a research direction of DT for 6G Network native AI. DTC is proposed as a digital virtual mapping of a wireless channel that reflects the entire process of channel fading states and variations in the physical world [179]. DTC is envisaged to continually interact with the wireless communications system channel requirements to support the system’s ability to adapt to complex communication environments actively. It can be applied in various typical 6G usage scenarios, offering significant performance enhancements in data rate, latency, spectral efficiency, energy efficiency, and link reliability [180]. To boost the capabilities of the DTC for the predictive 6G network, a novel DTC implementation platform is proposed, including data acquisition, environment sensing and reconstruction module to build the relationship between environment and channel for channel prediction, communication decision and interaction [181, 182]. A cluster-nuclei based channel model is proposed to analyze the mapping relationship between clusters and scatterers in the propagation environment, enabling channel reconstruction across various scenarios [183]. In addition, a radio environment knowledge (REK) pool is proposed to serve as a specific enabler for DTC. Based on the principles of REK pool, the electromagnetic wave property inspired REK construction method is proposed, which fills the gap in the interpretable mathematical representation of the relationship between adequate environment information and channel [184].
The ITU documentation defines the requirements and architecture of DT networks, emphasizing their core role as virtual representations of physical networks. By leveraging data collection, real-time interaction mapping, unified data models, and standardized interfaces, DTNs enable physical network analysis, diagnosis, simulation, and control. The ITU proposal introduces a “three-layer, three-domain, double closed-loop” architecture design, highlighting applications in complex network operations, optimization, innovation, and security strategy drills. DTNs’ functional requirements include data management, modeling, and lifecycle management, while service requirements focus on compatibility, scalability, and privacy, further supporting the intelligent, automated, and secure development of future networks.
DT technology is pivotal in developing 6G systems and various communication domains. By enabling precise modeling and real-time interaction with physical entities, DT provides practical solutions for optimizing resource allocation, enhancing communication security, and improving service management. For instance, DT optimizes computation offloading in industrial IoT in uRLLC links. Integrating DT with blockchain and FL technologies in wireless communication strengthens system security and privacy protection. In the vehicular network domain, DT leverages traffic prediction, edge computing, and load-balancing techniques to enhance real-time data processing and overall system efficiency. DT is driving innovation and development across multiple dimensions of communication systems, demonstrating its transformative potential. The application of DT technology in communication systems is currently experiencing rapid development, showcasing significant potential across various domains, particularly in the research and development of 5G and the emerging 6G networks. In industrial IoT environments, DT technology plays a crucial role. For instance, the authors of [185] address minimizing computation offloading delays in uRLLC links by leveraging DT-enabled wireless edge networks. The study focuses on modeling the computational capabilities of edge servers and optimizing system-wide resource allocation.
@@
To enable wireless communication systems based on proactive online learning, the authors of [186] introduce a DT design framework. This framework considers aspects such as twin object access, security and privacy, and air interface design, while also discussing the challenges and opportunities of DT in wireless systems. Additionally, the authors of [187] propose a blockchain-empowered FL framework to enhance communication security and data privacy within DT-enabled edge networks. The study incorporates asynchronous aggregation mechanisms and DT-driven RL to optimize user scheduling and spectrum resource allocation.
@@
DT technology has also seen applications in vehicular networks. The authors of [188] present a DT-assisted real-time traffic data prediction method capable of delivering accurate short-term forecasts for traffic flow and speed. The authors of [189] propose an innovative vehicular edge computing network combining DT and multi-agent learning. This approach helps uncover potential matches for edge services among large-scale vehicles, reducing the complexity of service management. Similarly, the authors of [190] introduce a DT-based load-balancing prediction model for autonomous vehicles, which accurately predicts road network conditions while ensuring high data transmission security. Furthermore, the authors of [191] explore the use of DT in vehicular edge computing, developing models that reflect the real-time state of the vehicular environment. The study introduces two metrics, Quantum DT and Cognitive DT, to evaluate the quality and cost of DT models for edge nodes, optimizing system performance and reducing costs. These studies highlight DT’s transformative potential across communication systems, from resource optimization in industrial IoT to service management and predictive modeling in vehicular networks.
The application of DT technology in communication systems has made initial progress but is still in the exploratory stage. In the future, DT technology needs to improve model construction methods and accuracy further, optimize real-time performance and latency, enhance security and privacy, and reduce resource consumption and costs. With AI and ML technology advancements, DT is expected to provide more intelligent network O&M solutions, achieving more efficient fault prediction, network optimization, and service quality management.
@@ 3.2.3 Computing Power Network and AI
The NET4AI architecture requires powerful computing service capabilities because wireless networks show great interest in computing tasks related to high-performance computing, such as ultra-large-scale data processing and DL. The computing power network is a key enabling technology for computing services[192]. It achieves ubiquitous computing interconnection through mutual sense and collaboration between the network and computing resources[193], promotes efficient harmony between cloud, edge and terminal resources, and serves all computing tasks in the network, especially AI-related computing services.
@@
The computing power network is a new type of information infrastructure. As shown in Figure 18, it supports computing power discovery, computing power management, task offloading and network computing, and based on this, it builds an Internet network composed of massive data, efficient computing power and ubiquitous intelligence, bringing computing services to every person, family and organization[194, 193]. Essentially, the computing power network is a service for computational resources. Its core idea is to connect geographically distributed computing power center nodes through new network technologies[192], dynamically sense the status of computing power resources in real time, then coordinate the allocation and scheduling of computing tasks, transmit data, and form a network that senses, allocates, and schedules computing power on a global scale. On this basis, computing power, data, and application resources are gathered and shared[195]. The technical principles of the computing power network can be divided into three aspects: 1) Real-time and accurate computing power discovery: Based on real-time network layer sense, computing power information (such as location and operating status) can be collected, whether it is concentrated in the cloud or distributed throughout the wireless network, thereby achieving rapid sense and discovery of computing power and advance routing calculation planning to ensure real-time and accurate provision of computing resources in the service. 2) Flexible and dynamic service scheduling: Dynamically plan and provide on-demand services based on the SLA requirements of the business. Through real-time resource status sense and accurate, flexible, and dynamic scheduling, business traffic can be quickly matched to the optimal node, achieving seamless service switching and delivery and enhancing user experience. 3) Consistency of user experience: The computing power network realizes the accurate sense and flexible allocation of computing resources between the cloud, edge and terminal devices, ensures a comprehensive and consistent user experience based on collaborative scheduling, and eliminates the need for users to care about the location or deployment status of computing resources within the network.
Currently, the computing power network is widely recognized in both academia and industry and primarily includes:
@@
Computational power measurement and modeling: This is a foundational technology for providing computational power services[196, 193]. In the future, computational power providers in computational power networks will not be confined to a specific data center or computing cluster. They will include ubiquitous computational power from cloud, edge, and end devices. Efficient sharing of this ubiquitous computational power through network connections requires accurate sensing of the computational capacity of these heterogeneous chips, the business types suitable for different chips, and their locations in the network, as well as effective management and supervision.
@@
Computational power routing based on resource awareness: In computational power networks, after measuring and modeling computational resources, the information is encoded and loaded into network control layer packets for sharing[197, 192]. The network control layer makes network decisions based on shared computational resource information, guiding business routing to different computational resource pools or through collaboration between computational resource pools for business processing, thus enabling network awareness of computational resources to guide global routing.
In-network computing (INC): Leveraging the deployment of programmable network technologies, INC processes packets within the network[198]. Sharing INC power using open and programmable heterogeneous resources accelerates data processing close to the source without altering the original business operating mode, reduces application response delays, and simplifies application deployment processes.
In summary, the computing power network provides stable and reliable computing resources for AI, using idle computing power in the network to complete the computing requirements of large-scale AI tasks, improving computing efficiency and reducing costs. Users do not need to build their computing infrastructure, enhancing the accessibility of AI computing and promoting the widespread application and adoption of AI technology. The computing power network integrates distributed computing resources to provide efficient, flexible, and scalable computing power to meet the real-time and environmental requirements of modern computing environments. It supports users to dynamically adjust resource scale according to demand to adapt to AI applications of different scales and complexities[197]. For example, users can request a large amount of idle computing power through the computing power network to support the high computing power requirements in new applications such as XR and AR.
@@
In addition, the computing power network also can be combined with other novel technologies. For example, by adopting blockchain technology[199], the computing power network can provide AI with a secure and privacy-protected computing environment, ensuring the transparency and traceability of the computing process. The introduction of technologies such as smart contracts can further protect user privacy and data security, allowing users to safely use the computing power provided by the computing power network to process sensitive data and tasks, thereby improving overall security and credibility. The composability of the computing power network gives it a brighter future in 6G networks.
@@
With the explosive growth of data volume in the network and computing requirements becoming increasingly complex, the computing power network has become the focus of attention of many researchers to solve these complex challenges. The authors of [200] implemented a Kubernete-based prototype testing platform for the computing power network using a microservices architecture, achieving key enabling technologies for the computing power network, including computational modeling, sense, announcement, and offloading. The authors of [201] described a lightweight multi-cluster hierarchical edge resource scheduling scheme based on a cloud-native resource scheduling mechanism, successfully managing and deploying many heterogeneous edge devices within a unified framework using a lightweight cloud-native platform. The authors of [195] proposed an integrated ICT network architecture of “connectivity + computing + intelligence,” integrating network, computing, storage, and applications into a unified design. The network can sense computational resources and perform related management and control. The authors of [202] designed an on-demand compute resource scheduling model and proposed a two-stage evolutionary search scheme. Numerical results indicate that this scheme significantly improves the efficiency and stability of task offloading and compute resource scheduling in edge computing networks. Detailed power analysis of several INC use cases in [198] showed that INC becomes more energy-efficient at very low processing loads, with increased processing loads having little impact on INC power consumption. The computing power network provides support and empowerment for AI in many ways. In [203], the fog computing network offers powerful computing power support for AI model training, reasoning and other tasks. This helps shorten model training cycles, improve training efficiency, and enable AI models to iterate and optimize faster. The study presented in [204] takes DNNs as a typical AI application, and formulates optimization problems under the constraints of energy consumption, delay, computing and communication resources to efficiently utilize computing resources and avoid resource waste.
@@ 3.2.4 Secure Multi-Party Computation
Ubiquitous interconnected services will generate vast amounts of user data, which face risks of privacy leakage and illegal usage during the processes of collection, storage, and transmission. In highly complex, heterogeneous, and dynamically evolving network environments, it is imperative to ensure data security while achieving efficient collaboration.
As network technology continues to evolve, so do network attack methods, becoming increasingly sophisticated. conventional security mechanisms are often inadequate to counter these emerging threats. As shown in Figure 19, secure multi-party computation (SMPC) [205] is an encryption protocol that allows multiple parties to jointly compute a function while keeping their input data private, . Each party obtains the correct computation result without knowing the other parties’ private inputs. SMPC has characteristics such as decentralization, input data security, and accurate computation results. It enables AI to handle complex network environments, protect user privacy and data security, support emerging application scenarios, and promote data sharing and utilization.
@@
(1) Garbled Circuits This technique involves compiling computational logic into a circuit and encrypting each gate (such as AND, XOR, etc.) within the circuit. Participants can complete the computation by interacting with the encrypted information without knowing the specific logic of the circuit.
@@
(2) Oblivious Transfer Oblivious Transfer is a protocol that allows one party (e.g., the sender) to send one of many pieces of information to another party (the receiver), where the receiver can choose only one piece of information. The sender does not know which piece of information the receiver has chosen.
@@
(3) Secret Sharing This method involves splitting a secret into multiple shares and distributing them to different participants. The original secret can only be reconstructed when a sufficient number of shares are combined.
@@
(4) Homomorphic Encryption Homomorphic Encryption is a special encryption method that allows specific computations (such as addition and multiplication) to be performed on encrypted data. The result of these computations, when decrypted, corresponds to the result of the same computations performed on the original data.
@@
SMPC, based on the aforementioned technologies, is widely applied in areas such as privacy-preserving ML. Recent research has leveraged SMPC to achieve secure model aggregation in FL, which involves training ML models on numerous decentralized devices while keeping the training data localized. This collaborative approach allows multiple parties to jointly train models without exposing their training data. For example, to counter reverse attacks on data, a privacy-preserving data aggregation mechanism based on secret sharing techniques in SMPC can efficiently protect FL against such attacks [211]. To address the issues of high communication overhead and poor scalability in conventional SMPC, a two-phase SMPC-based FL framework has been proposed [212]. This framework enables multiple clients to collaboratively train AI models while protecting their data privacy. To mitigate the high communication costs and poor scalability associated with aggregating global models at only some nodes, a hierarchical FL architecture using SMPC has been developed. This architecture improves communication efficiency and scalability without compromising privacy [213].
In the future mobile network, data is often characterized by high volume, high frequency, and asynchronous communication, presenting new challenges for the implementation of FL. A novel hybrid FL architecture has been proposed, which combines FL with Trusted Execution Environments, SMPC and Beidou satellites. This architecture achieves secure key distribution, encryption, and decryption, and provides verification mechanisms for each participant to ensure the security of local data [214]. SMPC can also be combined with blockchain technology to track and mitigate malicious behavior, achieving a secure and trustworthy collaborative edge learning framework [215].
SMPC can protect data privacy while enabling collaborative computation among multiple parties, making it of significant and widespread application in the 6G network. SMPC offers an effective solution to the contradiction between data privacy protection and collaborative computing. It plays a crucial role in scenarios such as privacy-preserving FL, distributed resource allocation and optimization, secure data fusion and analysis, distributed identity verification and key management, and anti-fraud billing systems. By leveraging encrypted computation, SMPC ensures the secure processing of sensitive information. In FL, SMPC enables multiple devices or nodes to train models without sharing raw data, widely applied in scenarios like smart healthcare and intelligent cities. In resource allocation and optimization, SMPC protects sensitive data from each party, facilitating efficient collaboration in tasks such as spectrum resource allocation and edge computing load balancing. In data fusion and intelligent analysis, SMPC ensures the privacy of cross-domain data statistics and modeling, such as user behavior analysis and network traffic prediction. In identity verification and key management, SMPC mitigates the risk of single-point failures in centralized schemes, supporting secure access for IoT devices and edge computing nodes. In dynamic billing systems, SMPC guarantees the privacy and fairness of user billing calculations, applied in international roaming and dynamic pricing models.
Furthermore, SMPC also plays a role in content distribution optimization, network security protection, and intelligent supply chain management. By protecting data privacy, it enables cross-domain collaboration, such as optimizing content transmission paths, collaborative threat detection, and supply chain inventory management. Combined with technologies like homomorphic encryption and secret sharing, SMPC can address the privacy challenges of data sharing in the 6G network while optimizing protocol performance to meet the demands of low latency and high bandwidth. This not only enhances the privacy and security of the network but also drives the intelligent and distributed collaborative development of the 6G network.
@@ 3.2.5 Semantic Communication
Semantic communication, commonly relying on DNNs, is considered as a promising technology in 6G. Comparing with the traditional syntactic communication that focuses on the accurate transmission of data in bits, semantic communication only transmits the semantics of the message rather than the whole raw data. In this way, the amounts of data to be transmitted are largely reduced. Meanwhile, it can understand the data context, preserving essential semantic relationships when transmitting information, thereby aiding downstream AI systems in making more intelligent and accurate decisions. These characteristics are to meet the demands of some new emerging data-hungry applications in 6G, e.g., holographic communication, XR, show in Figure 20.
Figure 20: Some semantic communication enabled data-hungry applications in 6G
Figure 21 provides a basic semantic communication model, encompassing three essential components: the semantic encoder, the channel, and the semantic decoder. At the transmitter, semantic encoder extracts and encodes SI of the source data. Specifically, this process involves extracting the SI and compressing or removing the irrelevant information. To achieve this, the raw data is encoded by a neural network, donated as ${\mathcal{F}}_{\theta}$, which outputs the semantics that retains the critical meaning while discarding non-essential details. The compact semantics are then transmitted over a noisy physical channel, such as a wireless channel. As such, noises or interferences are introduced, thereby resulting in the semantic distortion. At the receiver, the semantic decoder decodes the received data. Specifically, this process involves “understanding” and inferring the semantic content sent by the transmitter. To accomplish this, the semantic decoder, donated as ${\mathcal{G}}_{\phi}$, processes the received data with the goal of mapping it back to the source data. This ensures that the reconstructed meaning accurately reflects the source message, even in the presence of potential transmission distortion.
In addition, 6G networks are full of multi-modal data from heterogeneous sources, such as text, image, audio, point cloud and etc. Despite the heterogenous data structure, the semantics behind the multi-modal data may tightly correlated under some conditions. As such, the semantic correlation between the data in different modalities support the data redundancy reduction further.
@@
The authors of [216] provided a pioneering work in SI theory, establishing the mathematical framework for SI theory, which lays the foundation for semantic communication. The concept of modulation division multiple access technology was proposed in [217], which efficiently distinguishes users based on semantic features from the information dimension, thereby improving resource utilization in multi-user semantic communication systems, which contributes to multiuser semantic communication systems. Semantic extraction and recovery, also called semantic encoding and decoding, are the core components of semantic transceivers. Four dominant advanced semantic extraction techniques [218]: DL-based semantic extraction, RL-based semantic extraction, KB-assisted semantic extraction, and native semantic extraction. Moreover, most semantic communication studies concentrate on text transmission, audio transmission and recognition, image transmission and recognition, and video transmission and recognition from the perspective of data modality and task classification.
@@
Text Transmission: The authors of [219] developed a DL-based text semantic communication system named DeepSC, which performs joint semantic-channel coding to achieve superior performance gains compared to existing technologies, particularly at low SNRs. Recognizing that most semantic metrics are non-differentiable, a RL-based optimization paradigm for large-scale and complex text semantic transmission was introduced in [220]. This technique uses self-critical random iterative updates to train decoupled semantic transceivers, addressing the non-differentiable semantic channel optimization issue. The authors of [221] proposed a text-oriented semantic communication technique named semantic coding Reed Solomon hybrid automatic repeat request (HARQ) by combining HARQ and Reed-Solomon channel coding with semantic coding to tackle inefficiency and inflexibility with varying sentence lengths.
@@
Audio Transmission and Recognition: A DL-based audio semantic communication system named DeepSC-S was developed, as noted in [222, 223]. The audio semantic communication scheme employs a joint semantic encoder/decoder and channel encoder/decoder to mitigate the semantic distortion caused by the noises and interference from the wireless channels. Additionally, the authors of [224] explored FL-based audio semantic communication, developing a wav2vec-based autoencoder composed of CNNs. This system effectively encodes, transmits, and decodes audio SI, reducing communication overhead.
@@
Image Transmission and Recognition: The authors of [225] investigated a DeepJSCC image semantic communication architecture, where the encoder and decoder functions are parameterized by CNNs and jointly trained on the same dataset to minimize the mean squared error of the reconstructed images. Joint training enables DeepJSCC to avoid the cliff effect and promptly adapts to the varying SNR.
@@
Video Transmission and Recognition: The authors of [226] developed a variable-length DeepJSCC system that utilizes nonlinear transformations and conditional coding architectures to adaptively extract the semantic features of video frames. This system outperforms the traditional wireless video coding transmission on recognition accuracy.
@@
Semantic communication has unique advantages in semantic understanding, driving advancements in applications such as the tactile internet, holographic communication, XR, and human-machine symbiotic networks, fostering an increasingly intelligent and efficient society.

### v4

@@ 3.2.2 Digital Twins for 6G Network Native AI
In addition, the DT channel (DTC) has been proposed as a digital virtual mapping of a wireless channel that reflects the entire process of channel fading states and variations in the physical world [198]. DTC can be applied in various typical 6G usage scenarios, offering significant performance enhancements in data rate, latency, spectral efficiency, energy efficiency, and link reliability [199]. A novel DTC implementation platform was proposed, including data acquisition, environment sensing, and reconstruction module to build the relationship between environment and channel for channel prediction, communication decision, and interaction [200, 201]. A cluster-nuclei based channel model was proposed to analyze the mapping relationship between clusters and scatterers in the propagation environment, enabling channel reconstruction across various scenarios [202]. In addition, a radio environment knowledge (REK) pool was proposed to serve as a specific enabler for DTC. An electromagnetic wave property-inspired REK construction method was proposed, bridging the gap in the interpretable mathematical representation of the relationship between environmental information and the channel [203].
@@
The ITU documentation defines the requirements and architecture of DT networks DTNs[204], emphasizing their core role as virtual representations of physical networks. DTNs enable physical network analysis, diagnosis, simulation, and control. The ITU proposal introduces a “three-layer, three-domain, double closed-loop” architecture design, highlighting applications in complex network operations, optimization, innovation, and security strategy drills. For instance, DT optimizes computation offloading in industrial IoT in uRLLC links. Integrating DT with blockchain and FL technologies in wireless communication strengthens system security and privacy protection [205]. In industrial IoT environments, the authors of [206] minimized computation offloading delays in uRLLC links by leveraging DT-enabled wireless edge networks. To enable wireless communication systems based on proactive online learning, a DT design framework was introduced in [207]. This framework considers aspects such as twin object access, security and privacy, and air interface design. Additionally, in the work of [208], a blockchain-empowered FL framework was proposed to enhance communication security and data privacy within DT-enabled edge networks. The study incorporates asynchronous aggregation mechanisms and DT-driven RL to optimize user scheduling and spectrum resource allocation.
@@
An innovative vehicular edge computing network combining DT and multi-agent learning was proposed in [209], which uncovers potential matches for edge services among large-scale vehicles, reducing the complexity of service management. The authors of [210] presented a DT-assisted real-time traffic data prediction method capable of delivering accurate short-term forecasts for traffic flow and speed. A DT-based load-balancing prediction model for autonomous vehicles was introduced in [211], which accurately predicts road network conditions while ensuring high data transmission security. In [212], the authors explored the use of DT in vehicular edge computing, developing models that reflect the real-time state of the vehicular environment. The study introduced two metrics, Quantum DT and Cognitive DT, to evaluate the quality and cost of DT models for edge nodes.
@@ 3.2.3 Computing Power Network and AI
The NET4AI architecture requires powerful computing service capabilities because wireless networks show great interest in computing tasks related to high-performance computing, such as ultra-large-scale data processing and DL. The computing power network is a key enabling technology for computing services[213]. It achieves ubiquitous computing interconnection through mutual sense and collaboration between the network and computing resources[214], harmonizes cloud, edge, and terminal resources, and serves all computing tasks in the network, especially AI-related computing services.
@@
The computing power network is a new type of information infrastructure. As shown in Figure 16, it supports computing power discovery, computing power management, task offloading, and network computing[215, 214]. It connects geographically distributed computing power center nodes through new network technologies[213], and computing power, data, and application resources are gathered and shared[216].
The technical principles of the computing power network have three aspects: 1) Real-time and accurate computing power discovery; 2) Flexible and dynamic service scheduling; and 3) Consistency of user experience. Currently, the computing power network has gained wide recognition in both academia and industry and is classified into the following three categories.
@@
(1) Computational power measurement and modeling. This is a foundational technology for providing computational power services[217, 214]. In the future, computational power providers in computational power networks will not be confined to a specific data center or computing cluster. They will include ubiquitous computational power from cloud, edge, and end devices. Efficient sharing of this ubiquitous computational power through network connections requires accurate sensing of the computational capacity of these heterogeneous chips, the business types suitable for different chips, and their locations in the network, as well as effective management and supervision [218].
@@
(2) Computational power routing based on resource awareness. In computational power networks, after measuring and modeling computational resources, the information is encoded and loaded into network control layer packets for sharing[219, 213]. The network control layer makes network decisions based on shared computational resource information, guiding business routing to different computational resource pools or through collaboration between computational resource pools for business processing, thus enabling network awareness of computational resources to guide global routing.
@@
(3) In-network computing (INC). Leveraging the deployment of programmable network technologies, INC processes packets within the network[220]. Sharing INC power using open and programmable heterogeneous resources accelerates data processing close to the source without altering the original business operating mode, reduces application response delays, and simplifies application deployment processes. Computing power networks support users to dynamically adjust resource scale according to demand to adapt to AI applications of different scales and complexities[219]. In addition, the computing power network can also be combined with other novel technologies. For example, by adopting blockchain technology[221], the computing power network can provide AI with a secure and privacy-protected computing environment, ensuring the transparency and traceability of the computing process [222]. The introduction of technologies, such as smart contracts, can further protect user privacy and data security, allowing users to safely use the computing power provided by the computing power network to process sensitive data and tasks and improving overall security and credibility [223].
@@
In [224], a Kubernete-based prototype testing platform for the computing power network using a microservices architecture was implemented, achieving key enabling technologies for the computing power network, including computational modeling, sense, announcement, and offloading. The authors of [225] described a lightweight multi-cluster hierarchical edge resource scheduling scheme based on a cloud-native resource scheduling mechanism, successfully managing and deploying many heterogeneous edge devices within a unified framework using a lightweight cloud-native platform. An integrated ICT network architecture of “connectivity + computing + intelligence” was proposed in [216], which can sense computational resources and perform related management and control. Numerical results in [226] indicate that on-demand compute resource scheduling scheme significantly improves the efficiency and stability of task offloading and compute resource scheduling in edge computing networks. Detailed power analysis of several INC use cases in [220] showed that INC becomes more energy-efficient at very low processing loads, with increased processing loads having little impact on INC power consumption. In [227], the fog computing network offers powerful computing power support for AI model training, reasoning, and other tasks. This helps shorten model training cycles, improve training efficiency, and enable AI models to iterate and optimize faster. The study presented in [228] takes DNNs as a typical AI application, and formulates optimization problems under the constraints of energy consumption, delay, computing, and communication resources to efficiently utilize computing resources and avoid resource waste.
@@ 3.2.4 Secure Multi-Party Computation
In highly complex, heterogeneous, and dynamically evolving network environments, it is imperative to ensure data security while achieving efficient collaboration. As network technology continues to evolve, so do network attack methods, becoming increasingly sophisticated. Conventional security mechanisms are often inadequate to counter these emerging threats. As shown in Figure 17, secure multi-party computation (SMPC) [229] is an encryption protocol that allows multiple parties to jointly compute a function while keeping their input data private. Each party obtains the correct computation result without knowing the other parties’ private inputs. SMPC has characteristics such as decentralization, input data security, and accurate computation results. It enables AI to handle complex network environments, protect user privacy and data security, support emerging application scenarios, and promote data sharing and utilization.
@@
(1) Garbled Circuits technique involves compiling computational logic into a circuit and encrypting each gate (e.g., AND, XOR, etc.) within the circuit. Participants can complete the computation by interacting with the encrypted information without knowing the specific logic of the circuit.
@@
(2) Oblivious Transfer is a protocol that allows one party (e.g., the sender) to send one of many pieces of information to another party (the receiver), where the receiver can choose only one piece of information. The sender does not know which piece of information the receiver has chosen.
@@
(3) Secret Sharing involves splitting a secret into multiple shares and distributing them among participants. The original secret can only be reconstructed when a sufficient number of shares are combined.
@@
(4) Homomorphic Encryption is a special encryption method that allows specific computations (such as addition and multiplication) to be performed on encrypted data. The result of these computations, when decrypted, corresponds to the result of the same computations performed on the original data.
@@
Recent research has leveraged SMPC to achieve secure model aggregation in FL, which involves training ML models on decentralized devices while keeping the training data localized. To counter reverse attacks on data, a privacy-preserving data aggregation mechanism based on secret sharing techniques in SMPC can efficiently protect FL against such attacks [235]. To address high communication overhead and poor scalability in conventional SMPC, a two-phase SMPC-based FL framework has been proposed [236]. This framework enables multiple clients to collaboratively train AI models while protecting their data privacy. Furthermore, a hierarchical FL architecture utilizing SMPC has been developed to alleviate the communication costs and scalability issues associated with aggregating global models at only a few nodes, enhancing communication efficiency and scalability without compromising privacy [237]. A novel hybrid FL architecture has been proposed, which combines FL with Trusted Execution Environments, SMPC, and Beidou satellites. This architecture achieves secure key distribution, encryption, and decryption, and provides verification mechanisms for each participant to ensure the security of local data [238]. Additionally, SMPC can combine with blockchain technology to track and mitigate malicious behavior, achieving a secure and trustworthy collaborative edge learning framework [239].
SMPC protects data privacy while enabling collaborative computation across multiple parties, making it a valuable asset for 6G applications. It can contribute to FL, distributed resource allocation and optimization, secure data fusion and analysis, distributed identity verification and key management, and anti-fraud billing systems. By leveraging encrypted computation, SMPC ensures the secure processing of sensitive information. While SMPC enhances data sharing privacy and security in 6G networks, challenges remain, including computational and communication overhead, protocol standardization, and finding a balance between privacy protection and efficiency.
@@ 3.2.5 Semantic Communication
Semantic communication, commonly relying on DNNs, is considered a promising technology in 6G. Compared with traditional syntactic communication which focuses on the accurate transmission of data in bits, semantic communication only transmits the semantics of the message. It can also preserve essential semantic relationships when transmitting information, thereby aiding downstream AI tasks. These characteristics help meet the demands of new data-hungry 6G applications, e.g., holographic communication and XR, as shown in Figure 18.
Figure 18: Llustration of applications that require ultra-low-latency interaction in 6G networks, and deep learning-based semantic communications can greatly benefit these applications.
Figure 19 shows a basic semantic communication model, encompassing three essential components: the semantic encoder, the channel, and the semantic decoder [240]. At the transmitter, the semantic encoder extracts and encodes the SI of the source data. This process involves extracting the SI and compressing or removing the irrelevant information. To achieve this, the raw data is encoded by a neural network, denoted as ${\mathcal{F}}_{\theta}$, which outputs the semantics that retain the critical meaning while discarding non-essential details. The compact semantics are then transmitted over a noisy physical channel, such as a wireless channel. At the receiver, the semantic decoder decodes the received data. This process involves “understanding” and inferring the semantic content sent by the transmitter. To accomplish this, the semantic decoder, denoted as ${\mathcal{G}}_{\phi}$, processes the received data with the goal of mapping it back to the source data.
@@
The authors of [241] provided a pioneering work in SI theory, establishing the mathematical framework for SI theory. The concept of modulation division multiple access technology was proposed in [242], which efficiently distinguishes users based on semantic features from the information dimension, thereby improving resource utilization in multi-user semantic communication systems. Semantic extraction and recovery, also called semantic encoding and decoding, are the core components of semantic transceivers with four dominant advanced semantic extraction techniques [243]: DL-based semantic extraction, RL-based semantic extraction, KB-assisted semantic extraction, and native semantic extraction. Moreover, most semantic communication studies focus on text transmission, audio transmission, and recognition, image transmission and recognition, and video transmission and recognition from the perspective of data modality and task classification.
@@
(1) Text Transmission. The authors of [244] developed a DL-based text semantic communication system named DeepSC, which performs joint semantic-channel coding to achieve superior performance gains compared to existing technologies, particularly at low SNRs. Recognizing that most semantic metrics are non-differentiable, an RL-based optimization paradigm for large-scale and complex text semantic transmission was introduced in [245]. This technique uses self-critical random iterative updates to train decoupled semantic transceivers, addressing the non-differentiable semantic channel optimization issue. The authors of [246] proposed a text-oriented semantic communication technique named semantic coding Reed Solomon hybrid automatic repeat request (HARQ) by combining HARQ and Reed-Solomon channel coding with semantic coding to tackle inefficiency and inflexibility with varying sentence lengths.
@@
(2) Audio Transmission and Recognition. A DL-based audio semantic communication system named DeepSC-S was developed, as noted in [247, 248]. The audio semantic communication scheme employs a joint semantic encoder/decoder and channel encoder/decoder to mitigate the semantic distortion caused by the noises and interference from the wireless channels. The authors of [249] explored FL-based audio semantic communication, developing a wav2vec-based autoencoder composed of CNNs. This system effectively encodes, transmits, and decodes audio SI, reducing communication overhead.
@@
(3) Image Transmission and Recognition. The authors of [250] investigated a DeepJSCC image semantic communication architecture, where the encoder and decoder functions are parameterized by CNNs and jointly trained on the same dataset to minimize the mean squared error of the reconstructed images. Joint training enables DeepJSCC to avoid the cliff effect and adapt to the varying SNR.
@@
(4) Video Transmission and Recognition. The authors of [251] developed a variable-length DeepJSCC system that utilizes nonlinear transformations and conditional coding architectures to adaptively extract the semantic features of video frames. This system outperforms the traditional wireless video coding transmission on recognition accuracy.
@@
Semantic communication has unique advantages in semantic understanding, driving advancements in applications such as the tactile internet, holographic communication, XR, and human-machine symbiotic networks, fostering an increasingly intelligent and efficient network.